OpenStudy (anonymous):

Find the equation of the circle, in standard conic form, that has a diameter with endpoints A(-6,-4) and B(3,5). Your answer should be reduced and contain no decimals.

hero (hero):

\[(x + \frac{3}{2})^2 + (y - \frac{1}{2})^2 = \frac{81}{2}\]

OpenStudy (anonymous):

How did you get 40.5?

hero (hero):

I found the radius, then squared it

hero (hero):

but it says no decimals, so 81/2 instead

OpenStudy (anonymous):

Thanks. Got a couple more for you.

hero (hero):

I'm out. I have other obligations, sorry. Good luck

OpenStudy (anonymous):

Thanks for the help

hero (hero):

I was posting it because you left off the .5 r^2 = 40.5

hero (hero):

You'll see that on wolfram alpha, the original points are on the circle

hero (hero):

\[\text{approximate} \ne \text{reduce}\]

hero (hero):

reduce hints to reduce a fraction

hero (hero):

The question doesn't explicitly say to express h,k in the form of a fraction, instead it "hints" at it by using the word "reduce".

hero (hero):

Nah, I'm fine with you handwriting...as long as it is your own work and not some silly program. Once again, it is good to see you making the effort to post your own original solutions.